# Position Paper: What Can Large Language Models Tell Us about Time Series Analysis

### v1

@@ 4.3 Others
Beyond the previously discussed methods, another significant approach in temporal analysis involves building foundation models from scratch, as shown in Figure 3(c). This approach focuses on creating large, scalable models, both generic and domain-specific, aiming to emulate the scaling law (Kaplan et al., 2020) of LLMs. PreDcT (Das et al., 2023) used Google Trends data to build a vast time series corpus, altering a foundational model’s attention architecture for time prediction. Lag-Llama (Rasul et al., 2023) introduced a univariate probabilistic time series forecasting model, applying the smoothly broken power-law (Caballero et al., 2022) for model scaling analysis. TimeGPT (Garza & Mergenthaler-Canseco, 2023) further advanced time series forecasting, enabling zero-shot inference through extensive dataset knowledge. Recent efforts, including (Ekambaram et al., 2024), have focused on making these foundational models more efficient and applicable in specific areas like weather forecasting (Chen et al., 2023a), path planning (Sun et al., 2023), epidemic detection (Kamarthi & Prakash, 2023), and cloud operations (Woo et al., 2023).
@@ 4.4 Discussion
LLM-centric predictors have advanced significantly in time series analysis, outperforming most domain-specific models in few-shot and zero-shot scenarios. Tuning-based methods, with their adjustable parameters, generally show better performance and adaptability to specific domains. However, they are prone to catastrophic forgetting and involve high training costs due to parameter modification. While adapter layers have somewhat alleviated this issue, the challenge of expensive training persists. Conversely, non-tuning methods, offering text-based predictions, depend heavily on manual prompt engineering, and their prediction stability is not always reliable. Additionally, building foundational time series models from scratch involves balancing high development costs against their applicability. Therefore, further refinement is needed to address these challenges in LLM-centric predictors.
Our position: LLM-centric predictors, though burgeoning in time series analysis, are still in their infancy and warrant deeper consideration. Future advancements should not only build upon but also enhance time series foundation models. By harnessing unique LLM capabilities such as in-context learning and chain-of-thought reasoning, these advancements can overcome current limitations like catastrophic forgetting, and improve prediction stability and reliability.
@@ 5 LLM-empowered Agent for Time Series
In the subsequent section, we employ prompt engineering techniques to compel LLMs to assist in executing basic time series analytical tasks. Our demonstration reveals that LLMs undeniably possess the potential to function as time series agents. Nevertheless, their proficiency is constrained when it comes to comprehending intricate time series data, leading to the generation of hallucinatory outputs. Ultimately, we identify and discuss promising avenues that can empower us to develop more robust and reliable general-purpose time series agents.
@@ 5.1 Empirical Insights: LLMs as Time Series Analysts
This subsection presents experiments evaluating LLMs’ zero-shot capability in serving as agents for human interaction and time series data analysis. We utilize the HAR (Anguita et al., 2013) database, derived from recordings of 30 study participants engaged in activities of
@@
daily living (ADL) while carrying a waist-mounted smartphone equipped with inertial sensors. The end goal is to classify activities into four categories (Stand, Sit, Lay, Walk), with ten instances per class for evaluation. The prompts used for GPT-3.5 are illustrated in Figure 7, and the classification confusion matrix is presented in Figure 4. Our key findings include:
LLMs as Effective Agents: The experiments demonstrate that current LLMs serve adeptly as agents for human interaction and time series data analysis, producing accurate predictions as shown in Figure 7. Notably, all instances with label Stand were correctly classified, underscoring the LLMs’ proficiency in zero-shot tasks. The models exhibit a profound understanding of common-sense behaviors, encompassing various labels in time series classification, anomaly detection, and skillful application of data augmentation (Figure 8).
Interpretability and Truthfulness: Agent LLMs prioritize high interpretability and truthfulness, allowing users to inquire about the reasons behind their decisions with confidence. The intrinsic classification reasoning is articulated in natural language, fostering a user-friendly interaction.
Limitations in Understanding Complex Patterns: Despite their capabilities, current LLMs show limitations in comprehending complex time series patterns. When faced with complex queries, they may initially refuse to provide answers, citing the lack of access to detailed information about the underlying classification algorithm.
Bias and Task Preferences: LLMs display a bias towards the training language distributions, exhibiting a strong preference for specific tasks. In Figure 7, instances of Lay are consistently misclassified as Sit and Stand, with better performance observed for Sit and Stand.
(a) Align
@@
Hallucination Problem: The LLMs are susceptible to hallucination, generating reasonable but false answers. For instance, in Figure 8, augmented data is merely a copy of given instances, although the model knows how to apply data augmentation: These instances continue the hourly trend of oil temperature and power load features, maintaining the structure and characteristics of the provided dataset. Subsequent inquiries into the misclassification in Figure 4, particularly regarding why LLMs classify Lay instances as Sit and Stand, elicit seemingly plausible justifications (see Table 1). However, these justifications expose the model’s inclination to fabricate explanations.
@@ 5.2 Key Lessons for Advancing Time Series Agents
In light of the empirical insights from earlier experiments, it is apparent that LLMs, when serving as advanced time series analytical agents, exhibit notable limitations when dealing with questions about data distribution and specific features. Their responses often show a reliance on requesting additional information or highlight an inability to provide accurate justifications without access to the underlying model or specific data details.
To surmount such limitations and develop practical time series agents built upon LLMs, it becomes paramount to seamlessly integrate time series knowledge into LLMs. Drawing inspiration from studies that have successfully injected domain-specific knowledge into LLMs (Wang et al., 2023b; Liu et al., 2023b; Wu et al., 2023; Schick et al., 2023), we propose several research directions. These include innovative methods to enhance LLMs’ proficiency in time series analysis by endowing them with a deep understanding of temporal patterns and relevant contextual information.
  • Aligning Time Series Features with Language Model Representations (Figure 5a): Explicitly aligning time series features with pre-trained language model representations can potentially enhance the model’s understanding of temporal patterns. This alignment may involve mapping specific features to the corresponding linguistic elements within the model.
  • Fusing Text Embeddings and Time Series Features (Figure 5b): Exploring the fusion of text embeddings and time series features in a format optimized for LLMs is a promising avenue. This fusion aims to create a representation that leverages the strengths of LLMs in natural language processing while accommodating the intricacies of time series data.
  • Teaching LLMs to Utilize External Pre-trained Time Series Models (Figure 5c): The goal here is to instruct the LLM on identifying the appropriate pre-trained time series model from an external pool and guiding its usage based on user queries. The time series knowledge resides within this external model hub, while the LLM assumes the role of a high-level agent responsible for orchestrating their utilization and facilitating interaction with users.
Differentiating from approaches like model repurposing or fine-tuning on specific tasks, the focus of future research should be on harnessing the inherent zero-shot capabilities of LLMs for general pattern manipulation. Establishing a framework that facilitates seamless interaction between users and LLM agents for solving general time series problems through in-context learning is a promising direction.
@@ 5.3 Exploring Alternative Research Avenues
Hallucination, a recurring challenge documented in various foundational models (Zhou et al., 2023b; Rawte et al., 2023; Li et al., 2023a), proves to be a noteworthy concern when employing LLMs as agents for time series data analysis, as observed in our experiments. This phenomenon involves the generation of content that deviates from factual or accurate information. Addressing hallucination in this context commonly relies on two primary methods: the identification of reliable prompts (Vu et al., 2023; Madaan et al., 2023) and the fine-tuning of models using dependable instruction datasets (Tian et al., 2023; Zhang et al., 2023a). Nevertheless, it is crucial to recognize that these approaches often demand substantial human effort, introducing challenges related to scalability and efficiency. While certain initiatives have sought to integrate domain-specific knowledge into ICL prompts (Da et al., 2023a; Yang et al., 2022b) and construct instruction datasets tailored to specific domains (Liu et al., 2023b; Ge et al., 2023), the optimal format of instructions or prompts for effective time series analysis remains unclear. Exploring and establishing guidelines for crafting impactful instructions within the context of time series analysis represents a compelling avenue for future research.
Beyond this, there are ongoing concerns regarding the alignment of time series agents with human preferences (Lee et al., 2023), for instance, with a focus on ensuring the generation of content that is both helpful and harmless (Bai et al., 2022). These unresolved issues underscore the imperative necessity for the development of more robust and trustworthy time series agents. Moreover, the world is perpetually in a state of flux, and the internet undergoes continuous evolution, witnessing the addition of petabytes of new data on a daily basis (Wenzek et al., 2019). In the realm of time series analysis, the evolving pattern assumes greater significance due to the inherent concept drift in time series data (Tsymbal, 2004), where future data may exhibit patterns different from those observed in the past. Addressing this challenge involves enabling agents to continually acquire new knowledge (Garg et al., 2023) or adopting a lifelong learning pattern without the need for expensive retraining — a crucial aspect for time series agents.
Our Position: Current LLMs excel as agents for human interaction and time series data analysis, but they encounter issues such as occasional inaccuracies and a tendency toward hallucination. To improve their reliability in decision-making, it is crucial to develop guidelines for effective instructions and to incorporate domain-specific knowledge. Overcoming challenges like hallucination, aligning with human preferences, and adjusting to evolving time series data is key to maximizing their capabilities and minimizing risks. Our future vision is to develop robust and adaptable LLM-empowered agents that can adeptly handle the intricacies of time series analysis.
@@ 6 Further Discussion
Our perspectives serve as a starting point for ongoing discussion. We acknowledge that readers may have diverse views and be curious about aspects of LLM-centric time series analysis not addressed previously. Below, we objectively examine several of these alternate viewpoints:
Accountability and Transparency:
Despite intense public and academic interest, LLMs remain somewhat enigmatic, raising fundamental questions like their capabilities, operational mechanisms, and efficiency levels. These concerns are also pertinent in LLM-centric time series analysis, especially in recent studies using prompting methods without altering off-the-shelf LLMs, such as PromptCast (Xue & Salim, 2023). To foster transparent and accountable LLM-centric time series models, we advocate two key considerations. Firstly, a robust scientific process should aim to understand underlying mechanisms, as suggested in (Gruver et al., 2023). Secondly, establishing a transparent development and evaluation framework is necessary, one that the community can adopt for achieving better clarity — this includes consistent model reporting, benchmarking results, clear explanations of internal processes and outputs, and effective communication of model uncertainties (Liao & Vaughan, 2023).
Privacy and Security:
LLM-centric time series analysis introduces significant privacy and security challenges. Considering that much of the real-world time series data is confidential and sensitive, concerns about data leakage and misuse are paramount. LLMs are known to sometimes memorize segments of their training data, which may include private information (Peris et al., 2023). Consequently, developing and deploying LLM-centric time series models necessitates implementing privacy-preserving measures against threats like adversarial attacks and unauthorized data extraction. Formulating and adhering to ethical guidelines and regulatory frameworks is also important (Zhuo et al., 2023). These should specifically address the complex challenges posed by the use of LLMs in time series analysis, ensuring their responsible and secure application.
Environmental and Computational Costs:
The environmental and computational costs of LLM-centric time series analysis are subjects of considerable concern. Critics contend that the current benefits of these models do not outweigh the substantial resources needed for their functioning. In response, we argue that: (1) much of the existing research in this area is still in its infancy, offering substantial opportunities for optimization in tandem with LLM development; (2) exploring more efficient alignment and inference strategies is worthwhile. This is especially pertinent given the large context windows needed for handling tokenized high-precision numerical data.
@@ 7 Conclusion
This paper aims to draw the attention of researchers and practitioners to the potential of LLMs in advancing time series analysis and to underscore the importance of trust in these endeavors. Our position is that LLMs can serve as the central hub for understanding and advancing time series analysis, steering towards more universal intelligent systems for general-purpose analysis, whether as augmenters, predictors, or agents. To substantiate our positions, we have reviewed relevant literature, exploring and debating possible directions towards LLM-centric time series analysis to bridge existing gaps.
@@
Impact Statements
@@ Appendix A Literature Review
Insight Miner (Zhang et al., 2023f), AmicroN (Chatterjee et al., 2023),
(Yu et al., 2023b), (Yu et al., 2023c), (Fatouros et al., 2024), leaf, text width=0.53edge=output-black] ] [Model-Based Enhancer, middle, text width=0.16edge=output-black [IMU2CLIP (Moon et al., 2023), STLLM (Anonymous, 2024b), (Qiu et al., 2023b),
TrafficGPT (Zhang et al., 2023e), (Li et al., 2023b), (Yu et al., 2023a), (Qiu et al., 2023a), leaf, text width=0.53edge=output-black] ] ] [LLM-centered Predictor
@@
 (Spathis & Kawsar, 2023),  (Mirchandani et al., 2023),  (Zhang et al., 2023g), leaf, text width=0.53edge=output-black] ] [Others, middle, text width=0.16edge=output-black [Lag-Llama (Rasul et al., 2023), PreDcT (Das et al., 2023), CloudOps (Woo et al., 2023),
@@
(Section 5),middle,anchor=west,edge=output-black, text width=0.14[External Knowledge, middle, text width=0.16edge=output-black [GPT3-VQA (Yang et al., 2022b), PromptGAT (Da et al., 2023a),
Open-TI (Da et al., 2023b), Planner (Huang et al., 2022), Sociodojo (Anonymous, 2024a),
ProgPrompt (Singh et al., 2023), Visual ChatGPT (Wu et al., 2023), leaf, text width=0.53edge=output-black] ] [Adapt Target Modality, middle, text width=0.16edge=output-black [Toolformer (Schick et al., 2023), LLaVA (Liu et al., 2023b), PIXIU (Xie et al., 2023), leaf, text width=0.53edge=output-black] ] ] ]

### v2

@@ 4.3 Others
Beyond the previously discussed methods, another significant approach in temporal analysis involves building foundation models from scratch, as shown in Figure 4(c). This approach focuses on creating large, scalable models, both generic and domain-specific, aiming to emulate the scaling law (Kaplan et al., 2020) of LLMs. More details can be found in (Liang et al., 2024), as this branch of methods is somewhat peripheral to our primary positions in this paper.
@@ 4.4 Discussion
LLM-centric predictors have advanced significantly in time series analysis, often outperforming many domain-specific models in few-shot and zero-shot scenarios. Tuning-based methods, with their adjustable parameters, generally show better performance and adaptability to specific domains. However, they are prone to catastrophic forgetting and involve high training costs due to parameter modification. While adapter layers have somewhat alleviated this issue, the challenge of expensive training persists. Conversely, non-tuning methods, offering text-based predictions, depend heavily on manual prompt engineering, and their prediction stability is not always reliable. Additionally, building foundational time series models from scratch involves balancing high development costs against their applicability. Therefore, further refinement is needed to address these challenges in LLM-centric predictors.
Our position:  LLM-centric predictors, though burgeoning in time series analysis, are still in their infancy and warrant deeper consideration. Our position posits a crucial hypothesis that LLMs excel at processing time series tasks. Future advancements may not only build upon but also involve pre-trained time series models. By harnessing the unique capabilities of LLMs, these advancements can further reduce tuning costs and improve prediction stability and reliability.
@@ 5 LLM-empowered Agent for Time Series
In the subsequent section, we employ prompt engineering techniques to compel LLMs to assist in executing basic time series analytical tasks. Our demonstration reveals that LLMs undeniably possess the potential to function as time series agents. Nevertheless, their proficiency is constrained when it comes to comprehending intricate time series data, leading to the generation of hallucinatory outputs. Ultimately, we identify and discuss promising avenues that can empower us to develop more robust and reliable general-purpose single-agent and multi-agent time series systems.
@@ 5.1 Empirical Insights: LLMs as Time Series Analysts
This subsection presents experiments evaluating the LLM’s zero-shot capability as an agent for human interaction and time series data analysis. We utilize the HAR (Anguita et al., 2013) database, derived from recordings of 30 study participants engaged in activities of
@@
daily living (ADL) while carrying a waist-mounted smartphone equipped with inertial sensors. The end goal is to classify activities into four categories (Stand, Sit, Lay, Walk), with ten instances per class for evaluation. The prompts used for GPT-3.5 are illustrated in Figure 8, and the classification confusion matrix is presented in Figure 5. Our key observations include:
LLM as Effective Analytical Agent. The experiments demonstrate that the LLM serves adeptly as an agent for human interaction and time series data analysis, producing accurate predictions as shown in Figure 8. Notably, all instances with label Stand were correctly classified, underscoring the LLMs’ proficiency in zero-shot tasks. The models exhibit a profound understanding of common-sense behaviors, encompassing various labels in time series classification, anomaly detection, and skillful application of data augmentation (Figure 9).
Interpretability and Truthfulness. This single-agent system prioritizes high interpretability and truthfulness, allowing users to inquire about the reasons behind their decisions with confidence. The intrinsic classification reasoning is articulated in natural language, fostering a user-friendly interaction.
Limitations in Understanding Complex Patterns. Despite their capabilities, current LLMs show limitations in comprehending complex time series patterns. When faced with complex queries, they may initially refuse to provide answers, citing the lack of access to detailed information about the underlying classification algorithm.
Bias and Task Preferences. LLMs display a bias towards the training language distributions, exhibiting a strong preference for specific tasks. In Figure 8, instances of Lay are consistently misclassified as Sit and Stand, with better performance observed for Sit and Stand.
(a) Aligning
@@
Hallucination. LLMs are susceptible to hallucination problem, generating reasonable but false answers. For instance, in Figure 9, augmented data is merely a copy of given instances, although the model knows how to apply data augmentation: These instances continue the hourly trend of oil temperature and power load features, maintaining the structure and characteristics of the provided dataset. Subsequent inquiries into the misclassification in Figure 5, particularly regarding why LLMs classify Lay instances as Sit and Stand, elicit seemingly plausible justifications (see Table 1). However, these justifications expose the model’s inclination to fabricate explanations.
@@ 5.2 Key Lessons for Advancing Time Series Agents
General-purpose single-agent (and multi-agent) systems remain a missing piece in the puzzle of modern time series analysis. In light of the empirical insights from earlier experiments, it is apparent that LLMs, when serving as advanced single-agent time series systems, exhibit notable limitations when dealing with questions about data distribution and specific features. Their responses often show a reliance on requesting additional information or highlight an inability to provide accurate justifications without access to the underlying model or specific data details.
To surmount such limitations and develop practical time series agents built upon LLMs, it becomes paramount to seamlessly integrate time series knowledge into LLMs. Drawing inspiration from studies that have successfully injected domain-specific knowledge into LLMs (Wang et al., 2023a; Liu et al., 2023b; Wu et al., 2023; Schick et al., 2024), we propose several research directions. These include innovative methods to enhance LLM-based single-agent systems’ proficiency in time series analysis by endowing them with a deep understanding of temporal patterns and relevant contextual information.
  • Aligning Time Series Features with Language Model Representations (Figure 6a). Explicitly aligning time series features with pre-trained language model representations can potentially enhance the model’s understanding of temporal patterns. This alignment may involve mapping specific features to the corresponding linguistic elements within the model.
  • Fusing Text Embeddings and Time Series Features (Figure 6b). Exploring the fusion of text embeddings and time series features in a format optimized for LLMs is a promising avenue. This fusion aims to create a representation that leverages the strengths of LLMs in natural language processing while accommodating the intricacies of time series data.
  • Teaching LLMs to Utilize External Tools (Figure 6c). The goal here is to instruct the LLM to identify the appropriate pre-trained time series models or analytical tools from an external “toolbox” and guide their usage based on user queries. The time series knowledge resides within this external model hub, while the LLM assumes the role of a high-level agent, responsible for orchestrating their utilization and facilitating interaction with users.
Differentiating from approaches like model repurposing or fine-tuning on specific tasks, the focus of future research should be on harnessing the inherent zero-shot capabilities of LLMs for general pattern manipulation. Establishing a framework that facilitates seamless interaction between users and LLM agents for solving general time series problems through in-context learning is an exciting direction.
@@ 5.3 Exploring Alternative Research Avenues
Hallucination, a recurring challenge in various foundational models (Zhou et al., 2023b; Rawte et al., 2023; Li et al., 2023), is a significant concern in deploying LLM-based agent systems for time series analysis, as our experiments have shown. Addressing this issue typically involves two methods: identifying reliable prompts (Vu et al., 2023; Madaan et al., 2024) and fine-tuning models with dependable instruction datasets (Tian et al., 2023; Zhang et al., 2023a). However, these approaches require substantial human effort, posing scalability and efficiency challenges. Some initiatives integrate domain-specific knowledge into ICL prompts (Da et al., 2023; Yang et al., 2022b) and construct instruction datasets for specific domains (Liu et al., 2023b; Ge et al., 2023), but the best formats for instructions or prompts for effective time series analysis are still unclear. Developing guidelines for crafting impactful instructions in time series analysis is a promising area for future research.
Multi-agent time series systems is another promising direction composed of multiple interacting time series agents. These systems can leverage the strengths of individual agents, each specialized in different aspects of time series analysis, to provide more comprehensive and accurate results (Cheng et al., 2024). For instance, one agent might excel in identifying patterns and trends, while another focuses on anomaly detection or forecasting. The collaboration among these agents can lead to more robust solutions, particularly in complex scenarios where single-agent time series systems fall short.
Ongoing concerns about aligning time series agents with human preferences (Lee et al., 2023), such as generating helpful and harmless content (Bai et al., 2022), highlight the need for more robust and trustworthy agents. Additionally, the internet’s constant evolution, adding petabytes of new data daily (Wenzek et al., 2020), accentuates the importance of handling concept drift in time series data (Tsymbal, 2004), where future data may differ from past patterns. Addressing this challenge requires enabling agents to continually acquire new knowledge (Garg et al., 2023) or adopting lifelong learning without costly retraining.
Our Position: LLMs hold promise as agent systems for various time series applications, yet they encounter challenges such as occasional inaccuracies and hallucination. Enhancing their reliability requires effective instruction guidelines and domain-specific knowledge integration. Aligning with human preferences and adapting to evolving time series are crucial for maximizing their capabilities and minimizing risks. Our vision is to develop robust LLM-empowered agents capable of handling time series complexities, including exploring multi-agent systems for improved performance and reliability.
@@ 6 Further Discussion
Our perspectives initiate ongoing discussion. Acknowledging diverse views and potential curiosities regarding LLM-centric time series analysis, we objectively examine several alternate viewpoints:
Accountability and Transparency.
LLMs remain somewhat enigmatic, raising fundamental questions about their capabilities, mechanisms, and efficiency levels, especially in recent studies like PromptCast (Xue & Salim, 2023). We advocate for understanding underlying mechanisms (Gruver et al., 2023) and establishing transparent development and evaluation frameworks, including consistent model reporting and clear explanations of internal processes and outputs (Liao & Vaughan, 2024).
Privacy and Security.
LLM-centric time series analysis poses significant privacy and security challenges due to the sensitivity of industrial time series data. LLMs are known to sometimes memorize segments of their training data, which may include private information (Peris et al., 2023). Measures against threats like data leakage and misuse are crucial, along with ethical guidelines and regulatory frameworks to ensure responsible and secure application (Zhuo et al., 2023).
Environmental and Computational Costs.
Critics highlight the environmental and computational costs of LLM-centric time series analysis, suggesting optimization opportunities in LLM development and exploring more efficient alignment and inference strategies, especially for handling tokenized high-precision numerical data.
@@ 7 Conclusion
This paper aims to draw the attention of researchers and practitioners to the potential of LLMs in advancing time series analysis and to underscore the importance of trust in these endeavors. Our key position is that LLMs can serve as the central hub for understanding and advancing time series analysis, steering towards more universal intelligent systems for general-purpose analysis, whether as enhancers, predictors, or agents. To substantiate our positions, we have reviewed relevant literature, exploring and debating possible directions towards LLM-centric time series analysis to bridge existing gaps.
@@
Acknowledgements
This material is based on research partially sponsored by the CSIRO – National Science Foundation (US) AI Research Collaboration Program. S. Pan was supported in part by the Australian Research Council (ARC) under grants FT210100097 and DP240101547. Y. Liang was supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620).
Impact Statement
@@ Appendix A Literature Review
Insight Miner (Zhang et al., 2023c), AmicroN (Chatterjee et al., 2023),
(Yu et al., 2023b), (Yu et al., 2023c), (Fatouros et al., 2024), leaf, text width=0.53edge=output-black] ] [Model-based Enhancer, middle, text width=0.16edge=output-black [IMU2CLIP (Moon et al., 2023), STLLM (Zhang et al., 2024c), (Qiu et al., 2023b),
TrafficGPT (Zhang et al., 2024d), (Li et al., 2024), (Yu et al., 2023a), (Qiu et al., 2023a), leaf, text width=0.53edge=output-black] ] ] [LLM-centered Predictor
@@
 (Spathis & Kawsar, 2023),  (Mirchandani et al., 2023),  (Zhang et al., 2023d), leaf, text width=0.53edge=output-black] ] [Others, middle, text width=0.16edge=output-black [Lag-Llama (Rasul et al., 2023), PreDcT (Das et al., 2024), CloudOps (Woo et al., 2023),
@@
(Section 5),middle,anchor=west,edge=output-black, text width=0.14[External Knowledge, middle, text width=0.16edge=output-black [GPT3-VQA (Yang et al., 2022b), PromptGAT (Da et al., 2023),
Open-TI (Da et al., 2024), Planner (Huang et al., 2022), Sociodojo (Cheng & Chin, 2024),
ProgPrompt (Singh et al., 2023), Visual ChatGPT (Wu et al., 2023), leaf, text width=0.53edge=output-black] ] [Adapt Target Modality, middle, text width=0.16edge=output-black [Toolformer (Schick et al., 2024), LLaVA (Liu et al., 2023b), PIXIU (Xie et al., 2023), leaf, text width=0.53edge=output-black] ] ] ]